# Distributed acoustic sensing for ocean applications

### v2

@@ VI.2 Ocean surface waves, nearshore processes, and tsunamis
Ocean-bottom DAS is sensitive to cable strain induced by the pressure perturbations from surface gravity waves (SGW; Fig. 19).lindsey2019illuminating ; sladen2019distributed ; williams2019distributed Recent work has demonstrated that DAS can quantify sea surface statistics, such as significant wave height and average wave period. glover2024comparisons ; glover2024measuring ; meule2024reconstruction ; smith2023observations
Water depth and gauge length determine whether SGWs can be observed by ocean-bottom DAS. All SGWs follow the dispersion relation
@@
where $\omega$ is angular frequency, $k$ is wavenumber, $g$ is gravitational acceleration, and $h$ is water depth (Fig. 19b). Wind waves and swell in the 0.05–0.5 Hz band therefore have wavelengths of 6–600 m in deep water ($kh\gg 2\pi$), decreasing as the waves shoal into shallow water. Choice of appropriately short gauge length and pulse width (19) is therefore important for resolving the higher-frequency components of the SGW spectrum. glover2024comparisons
The pressure field of SGWs decays exponentially with depth as a function of the wavelength, such that the ratio of the pressure at the seafloor to the pressure at the sea surface ($p_{0}=\rho gA$, where $A$ is wave height) is
$\frac{p}{p_{0}}=\frac{1}{\cosh{kh}}$ (34)
@@
where $h$ is water depth and $k=k(\omega)$ is given by (33). In shallow water ($kh\ll 2\pi$), this hydrodynamic transfer function is approximately constant, but differs by many orders of magnitude in deeper water. Consequently, DAS can only observe wind waves and swell ($\sim 50-500$ mHz) in shallow water near the coast. Infragravity waves ($\sim 3-30$ mHz) and tsunamis ($\lesssim 20$ mHz) are observable at all depths.bromirski2002near ; taweesintananon2023distributed ; glover2024measuring ; shi2024multiplexed
@@
Figure 19: (color online) (a) Ocean surface gravity waves observed between channels 3000–4000 on the OOI RCA North cable from dataset A (depth ca. 50 m), band-pass filtered 0.05–0.5 Hz. (b) Frequency-wavenumber spectrum of the same data segment, showing propagation towards the shore at positive wavenumbers and the theoretical SGW dispersion relation (white line).
Comparison of DAS spectra with collocated buoys and pressure sensors in four recent studiesglover2024comparisons ; glover2024measuring ; meule2024reconstruction ; smith2023observations has shown that the DAS strain measurements are linearly correlated with SGW bottom pressure over a wide variety of sea states, cable types, and installation conditions. For a cable uniformly coupled to elastic sediment, the axial strain in the cable is identical to the axial strain in the surrounding sediment,kuvshinov2016interaction ; williams2024toward suggesting that the DAS sensitivity to SGW comes from seafloor compliance, as typically seen on ocean-bottom seismometers.crawford1991seafloor For DAS on unburied cables or where the sediment is nearly fluid, sensitivity to SGW should arise from direct pressure loading on the cable through the Poisson effect, which is a function of the cable construction. glover2024measuring The sensitivity could then range between $10^{-11}$ to $10^{-8}$ strain/Pa for buried cablesdoran2016infragravity and $10^{-12}$ to $10^{-10}$ strain/Pa for unburied cables.budiansky1979pressure These expectations are generally consistent with observationally derived transfer functions, though variability in sensitivity up to 6 orders of magnitude from cable to cable remains a challenge.glover2024comparisons ; glover2024measuring ; meule2024reconstruction ; smith2023observations
DAS is inherently multi-scale, with a spatial resolution span 1–105 m. Because cables are often buried in shallow water, high-fidelity wave measurements can be made in a continuous transect from the beach to the outer shelf and are immune to strong currents or sea ice. For example, a 36-km cable across the Beaufort Shelf measured the attenuation law of SGW in sea ice and mapped its spatial variability under changing ice conditions.smith2023observations In another study, array processing was applied to separate incident and shore-reflected SGWs to assess the influence of coastal reflection on microseism generation in the Mediterranean.guerin2022quantifying DAS array measurements of SGW dispersion in shallow water have also been used to study wave-current interaction and estimate depth-averaged current speed across meter to kilometer scales.williams2019distributed ; williams2022surface ; lin2024monitoring Recent observation of infragravity waves and small tsunamisshi2024multiplexed ; tonegawa2024high ; xiao2024detection also points to a future where the dense coastal network of seafloor communication and power cables can be leveraged to increase dramatically the density of the existing sparse buoy network for real-time tsunami detection and early warning.
@@ VI.3 Temperature sensing
DAS can directly measure changes in fiber temperature. Commonly, DAS refers to the family of technologies that measure phase changes of Rayleigh-scattered light due to variations of the optical path length due to external stimuli, regardless of whether these stimuli originate from mechanical vibrations or temperature fluctuations. The main distributed temperature sensing (DTS) is a distinct technology based on Raman scattering, see Fig. 1. The temperature sensitivity of DAS has been exploited in borehole geophysics to monitor fluid flow, where it is sometimes called distributed temperature gradient sensing.sidenko2022experimental ; titov_spe ; karrenbach2019fiber ; bradley2024
@@
For DAS, similar to (18), the phase change over a gauge length $L_{g}$ is a function of changes in the optical path length ($\Delta L/L_{g}$) and refractive index ($\Delta n/n$), which are modified by temperature through the thermo-elastic and thermo-optic effects
$\Delta\Phi$ $=\frac{4\pi nL_{g}}{\lambda}\left(\frac{\Delta L}{L_{g}}+\frac{\Delta n}{n}\right)$
$=\frac{4\pi nL_{g}}{\lambda}\left(\alpha_{T}+\psi\right)\Delta T,$ (35)
@@
where $\Delta T$ is a temperature change, $\alpha_{T}=\Delta L/(L_{g}\Delta T)$ is the coefficient of thermal expansion, and $\psi=\Delta n/(n\Delta T)$ is the thermo-optic coefficient, reflecting the relative change in the index of refraction with temperature.sidenko2022experimental Substituting typical values of these coefficients for silica fiber,hocker1979fiber the sensitivity to temperature is $\Delta\Phi/\Phi=7\times 10^{-6}$ 1/C, where $\Phi=4\pi nL_{g}/\lambda$. Combining this with the strain sensitivity (18), a temperature change of 1C induces the same phase change $\Delta\Phi$ as a mechanical strain of $10^{-5}$ m/m.zumberge2018measuring
@@
Internal gravity waves and turbulence in the continuously stratified ocean cause temperature oscillations at any fixed sensor. DAS recordings from seafloor cables offshore Japanide2021very , in the Canary Islandswilliams2023fiber , and Mediterranean Seapelaez2023high have revealed low-frequency transients over time-scales from minutes to weeks, inferred to represent temperature oscillations from internal tides and turbulence. These DAS temperature observations are qualitatively consistent with observations from Raman-based DTS deployed on seafloor cables.sinnett2020distributed ; davis2020fate
For oceanographic studies, the greatest advantage of Rayleigh-based DAS over Raman-based DTS is its range, which is typically 4 to 10 times longer than commercial DTS systems, such that a DAS interrogator installed on land can monitor temperature dynamics from the shallow shelf to the abyssal ocean. Compared to DTS, DAS also has improved sensitivity to small (milli-Kelvin scale) temperature fluctuations. However, DAS can only measure relative temperature, whereas Raman-based DTS can be calibrated to measure the absolute temperature at any point along a short ($<10-30$ km) cable.sidenko2022experimental ; williams2023fiber ; sinnett2020distributed
Because DAS is a single-component sensor, temperature and strain contributions to the measured phase cannot be disentangled. In some applications, temperature transients may be distinguishable from mechanical vibrations using array analysis, as performed above for acoustic signals (e.g., Fig. 8), because internal waves and turbulence occupy characteristic space-time scales.pelaez2023high However, given the gauge lengths of DAS (1–100 m), the smallest-scale temperature oscillations associated with turbulent eddies may be resolved in time but not in space, possibly contributing to the measurement noise in seismo-acoustic DAS. Cable burial insulates DAS from temperature transients, and temperature effects can be ignored at seismic and acoustic frequencies in buried cables.williams2023fiber If bespoke cables are used, DAS can be designed to orthogonalize strain and temperature by probing multiple fibers with different thermal properties.zumberge2018measuring
@@ VII Conclusions
This review provides an encompassing description of DAS technology and demonstrates its potential applications to ocean acoustics. We detail how DAS converts a fiber-optic cable into a distributed sensor of vibrational fields, such as propagating sound, substantiating that active optic sensing can be used as a proxy for passive acoustic monitoring of the environment. Based on the mathematical derivation of the underlying physics from electromagnetic to mechanical and acoustic quantities in DAS measurements, we demonstrate the effect of typical DAS acquisition parameters in signal processing. The potential of DAS technology for underwater acoustic applications, such as sound source detection, is demonstrated with the Ocean Observatories Initiative data.
Deriving the directional sensitivity of the distributed sensing modality in terms of conventional array signal processing, we explain that a limitation of current DAS systems in observing high-frequency signals stems from competing data acquisition requirements. Specifically, the maximum detectable frequency is determined by the repetition frequency of the pulse that interrogates the optical fiber, which, in turn, is inversely proportional to the measurement range. Additionally, the longer the transmitted pulse length, the higher the signal-to-noise ratio of the measurand but the less sensitive the system becomes to high-frequency vibrational fields.
@@ Acknowledgements.
This work was funded by the NATO Allied Command Transformation and by the Office of Naval Research under Grant N00014-21-1-2267.
@@ Appendix A Coherent optical reflectometry
The backscattered light at the optical fiber interrogator is detected with coherent optical reflectometry. In principle, coherent optical reflectometry mixes the backscattered signal returning from the interrogated fiber $E_{b}$ with a reference signal $E_{\text{ref}}$. The resulting signal is fed to the photodetector such that the output current $I_{pc}$ is proportional to the power of the resulting interferometric field He2021
@@
$=\lvert E_{\text{ref}}\rvert_{2}^{2}+\lvert E_{b}\rvert_{2}^{2}+\Re\left\{2E_{\text{ref}}E^{*}_{b}\right\}.$
@@
The reference signal $E_{\text{ref}}$ is, usually, a replica of the transmitted pulse $E_{p}$, potentially amplified and frequency shifted by an optical local oscillator
$\widetilde{E}_{\text{ref}}(\tau,\omega+\Delta\omega)=G_{\text{LO}}e^{j\left(\Delta\omega\tau+\phi(\tau)\right)}\widetilde{E}_{p}(\tau,\omega),$ (37)
@@
where $G_{\text{LO}}$ is the gain and $\Delta\omega$ the frequency shift introduced by the local oscillator, whereas $\phi(\tau)$ accounts for the time-dependent phase introduced by differences in the optical path between the reference and backscattered signal. Deriving from Eqs. (37), (12), and (14), the interference product becomes
$\widetilde{E}_{\text{ref}}(\tau,\omega\!+\!\Delta\omega)\widetilde{E}^{*}_{b}(\tau,\omega)=$ (38)
$\quad G_{\text{LO}}e^{j\left(\Delta\omega\tau+\phi(\tau)\right)}p\left(\frac{\tau}{T_{p}}\right)e^{j(\omega\tau+\phi_{0})}\cdot$
$\int\limits_{-\infty}^{\infty}\!\!p\left(\frac{\tau\!-\!2\frac{z_{s}}{c_{n}}}{T_{p}}\right)\!e^{-j\left(\omega\left(\tau\!-\!2\frac{z_{s}}{c_{n}}\right)+\phi_{0}\right)}s(z_{s})\mathrm{d}z_{s}=$
@@
$G_{\text{LO}}e^{j\left(\Delta\omega\tau+\phi(\tau)\right)}\!\!\!\!\!\int\limits_{-L_{p}/2}^{L_{p}/2}\!\!\!\!\!\!w(z)w(z\!-\!z_{s})s(z_{s})e^{j2kz_{s}}\mathrm{d}z_{s}=$
$G_{\text{LO}}e^{j\left(\Delta\omega\tau+\phi(\tau)\right)}\left[\widetilde{W}(k)\widetilde{W}^{*}(k)\widetilde{S}^{*}(k)\right].$
@@
Lowpass filtering the receiver output excludes the high-frequency signals around the nominal frequency $\omega$ and retains the signal at the difference frequency $\Delta\omega$
$I_{pc}(\tau)\propto\Re\left\{\widetilde{E}_{\text{ref}}(\tau,\omega\!+\!\Delta\omega)\widetilde{E}^{*}_{b}(\tau,\omega)\right\}=$ (39)
$\quad G_{\text{LO}}\cos\left(\Delta\omega\tau\!+\!\phi(\tau)\right)\Re\left\{\widetilde{W}(k)\widetilde{W}^{*}(k)\widetilde{S}^{*}(k)\right\}.$
@@
Couplers can be used to extract the phase of the detected signal [Hartog2017, , Sec. 3.3.1.1]. For example, an in-phase/quadrature (I/Q) demodulation systemwang2016coherent uses a $2\times 2$ channel coupler to output both the in-phase signal (39) and the corresponding quadrature signal by phase shifting the in-phase signal by $\pi/2$
$I_{pc}(\tau,\pi/2)\propto G_{\text{LO}}\sin\left(\Delta\omega\tau+\phi(\tau)\right)\Re\left\{\widetilde{W}(k)\widetilde{W}^{*}(k)\widetilde{S}^{*}(k)\right\}.$ (40)
Then, the phase of the detected signal can be calculated as
$\Delta\omega\tau+\phi(\tau)=\arctan\left(\frac{I_{pc}(\tau,\pi/2)}{I_{pc}(\tau)}\right)$ (41)
$\phi(\tau)=\arctan\left(\frac{\sin\left(\Delta\omega\tau+\phi(\tau)\right)}{\cos\left(\Delta\omega\tau+\phi(\tau)\right)}\right)-\Delta\omega\tau$
Since the backscattered field is a random process the output of the photodetector varies randomly, exhibiting a speckle-like behavior. However, averaging the signal at the photodetector over several transmissions the resulting output is proportional to the power spectral density of the pulse’s apodization function $\widetilde{C}_{ww}=<\widetilde{W}(k)\widetilde{W}^{*}(k)>_{p}$
$<\!I_{pc}\!>_{p}\propto G_{\text{LO}}e^{j\Delta\omega\tau}\Re\!\left\{\!<\!\widetilde{W}(k)\widetilde{W}^{*}(k)\widetilde{S}^{*}(k)\!>_{p}\!\right\}\!.$ (42)
Narrowband pulses result in a long autocorrelation response, whereas broadband excitation, e.g. frequency-modulated pulses, results in a shorter autocorrelation function and hence improves temporal resolution or, equivalently, spatial resolution. Figure 20 illustrates that the temporal resolution of a narrowband pulse is proportional to the pulse duration (a)–(c), whereas the resolution of a broadband pulse is inversely proportional to the bandwidth (d)–(f). Long pulses are required to achieve a sufficient signal-to-noise ratio (SNR). Broadband frequency modulated pulses (chirps) achieve effective pulse compression, hence finer resolution for the same pulse duration and consequently higher SNR.
Figure 20: Transmitted laser pulse characteristics. (a) Sinusoidal pulse of duration T multiplied by a tapered cosine window, (b) its frequency spectrum, and (c) autocorrelation function. (d) Linear frequency modulated pulse of duration T and bandwidth $\Delta f$ multiplied by a tapered cosine window, (e) its frequency spectrum, and (f) autocorrelation function.
Coherent optical time-domain reflectometry (COTDR) is used with narrowband sources and increases the SNR to improve detection by introducing a substantial gain to the local oscillator. Coherent optical frequency-domain reflectometry (COFDR) is used with broadband sources producing frequency-modulated pulses to increase the SNR by increasing the pulse length without compromising the resolution.

### v3

@@ VI.2 Ocean surface waves, nearshore processes, and tsunamis
Very low frequency (<<0.5 Hz) pressure perturbations from ocean surface gravity waves (SGW) also induce strain in a fiber-optic cable deployed on the seafloor, which can be measured with DAS lindsey2019illuminating ; sladen2019distributed ; williams2019distributed and used to quantify sea surface characteristics, such as significant wave height and average wave period. glover2024comparisons ; glover2024measuring ; meule2024reconstruction ; smith2023observations Figure 20 shows DAS observations of SGW along the north cable of the OOI Regional Cabled Array and their dispersion characteristics in the f-k spectra.
The dispersion relation for SGW [shown with a dotted line in Fig. 20(b)] varies with water depth $h$ as
@@
where $\omega=2\pi f$ is angular frequency at frequency $f$, $k=2\pi/\lambda$ is wavenumber at wavelength $\lambda$, and $g$ is gravitational acceleration. In deep water, i.e., for $h\geq\lambda/(2\pi)$, $\tanh(kh)\approx 1$ and $\lambda=g/(2\pi f^{2})$ from Eq. (38), hence wind waves and swell with frequencies around 0.05–0.5 Hz have wavelengths around 600–6 m. In shallow water, i.e., for $h\ll\lambda/(2\pi)$, $\tanh\left(kh\right)\approx kh$ and $\lambda=\sqrt{gh}/f$, hence the wavelength scales as $\sqrt{h}$, asymptotically approaching zero at the beach. The choice of an appropriately short gauge length and pulse width (see Eq. (25)) is, therefore, essential to resolve the smaller wavelengths in shallow water and accurately estimate parameters like significant wave height.glover2024comparisons
For a plane harmonic SGW with amplitude $A$ at the sea surface, the ratio of the pressure at the seafloor $p$ to the pressure at the sea surface $p_{0}=\rho gA$ is bromirski2002near
$\frac{p}{p_{0}}=\frac{1}{\cosh(kh)},$ (39)
where $k$ follows the dispersion relation in Eq. (38). In shallow water ($h\ll\lambda/(2\pi)$), $\cosh(kh)\approx 1$, hence the pressure $p=p_{0}$ does not depend on water depth. Conversely, in deep water ($h\geq\lambda/(2\pi)$), $\cosh(kh)\approx e^{kh}/2$, hence the pressure $p=p_{0}e^{-kh}/2$ decays exponentially with depth. For example, at 1 m water depth, a 0.02 Hz and a 0.2 Hz wave exert the same pressure at the seafloor, whereas, at 100 m water depth, the pressure due to a 0.02 Hz wave is more than $10^{7}$ times larger than the pressure due to a 0.2 Hz wave.taweesintananon2023distributed
@@
Consequently, water depth acts as a low-pass filter for the SGW spectrum, so DAS can only observe wind waves and swell ($\sim$0.05–0.5 Hz) in shallow water near the coast where the seafloor pressure is finite. Infragravity waves ($\sim$ 0.003–0.03 Hz) and tsunamis ($\lesssim 0.02$ Hz) are generally observable at all depths.bromirski2002near ; taweesintananon2023distributed ; glover2024measuring ; shi2024multiplexed
@@
Figure 20: (color online) (a) Ocean surface gravity waves observed between channels 3000–4000 (depth $\sim$50 m) on the OOI Regional Cabled Array north cable, band-pass filtered 0.05–0.5 Hz. The data are concatenated from four successive files starting from the file of Dataset A. (b) Frequency-wavenumber spectrum of the same data segment, showing propagation towards the shore at positive wavenumbers and the theoretical dispersion relation of surface gravity waves (white dotted line).
DAS enables wave measurement in a continuous transect from the beach to the outer shelf, and, since cables are often buried in shallow water, DAS is generally immune to strong currents or sea ice. DAS has been used to measure the attenuation of SGW in sea icesmith2023observations and track tsunamis generated by glacial calving events.graeff2024shaken Array processing of DAS data has been used to assess the influence of coastal reflection of SGWs on microseism generationguerin2022quantifying and quantify wave-current interaction in shallow water.williams2019distributed ; williams2022surface ; lin2024monitoring Observation of infragravity waves and small tsunamis with DASshi2024multiplexed ; tonegawa2024high ; xiao2024detection suggests that the existing global subsea network of fiber-optic cables can be used for real-time tsunami detection and early warning.
@@ VI.3 Temperature sensing
In addition to mechanical strain, DAS can directly measure temperature changes. For a temperature change $\Delta T$, the associated phase change over a gauge length $L_{g}$ is a function of changes in the optical path length ($\Delta L/L_{g}$) and refractive index ($\Delta n/n$), similarly to Eq. (23) [hocker1979fiber, , Eq. (16)],
$\Delta\Phi$ $=\frac{4\pi nL_{g}}{\lambda_{p}}\left(\frac{\Delta L}{L_{g}}+\frac{\Delta n}{n}\right)$
$\approx\frac{4\pi nL_{g}}{\lambda_{p}}\left(\alpha_{T}+\psi\right)\Delta T,$ (40)
@@
where $\alpha_{T}=L_{g}^{-1}\partial L/\partial T$ represents the thermo-elastic effect (expansion of the fiber with increasing temperature) and $\psi=n^{-1}\partial n/\partial T$ represents the thermo-optic effect (slower propagation of light with increasing temperature).hocker1979fiber ; sidenko2022experimental Substituting typical values of these coefficients for silica fiber,hocker1979fiber the sensitivity to temperature is $\Delta\Phi/\Phi=7\times 10^{-6}$ per 1 C temperature change, where $\Phi=4\pi nL_{g}/\lambda_{p}$. Comparing this with the strain sensitivity in Eq. (23), a temperature change of 1C induces the same phase change $\Delta\Phi$ as a mechanical strain of $10^{-5}$ m/m.zumberge2018measuring
Equation (40) indicates that DAS can be also used for distributed measurements of temperature changes. More common temperature sensing methods, i.e, distributed temperature sensing (DTS) based on Raman scattering and distributed strain and temperature sensing (DSTS) based on Brillouin scattering (Fig. 1), measure the amplitude rather than the phase of backscattered light.Hartog2017 Consequently, the sensitivity of both DTS and DSTS systems decreases with the measurement range, which can only be partly mitigated by increasing the temporal or spatial averaging.selker2006distributed One advantage of DAS is its long measurement range;koyamada2009fiber ; williams2023fiber see Sec. V.2.
DAS is more sensitive to very small (milliKelvin-scale) temperature fluctuations than DTS or DSTS.koyamada2009fiber ; seabrook2022comparison However, DAS can only measure relative temperature, whereas DTS can be calibrated to measure the absolute temperature.sidenko2022experimental ; sinnett2020distributed DTS only measures temperature and is unaffected by strain. In principle, DSTS can independently measure both strain and temperature by combining information from the Brillouin frequency shift and scattering power.clement2021b By contrast, temperature and strain contributions in DAS measurements cannot be disentangled, unless probing simultaneously multiple fibers with different thermal properties.zumberge2018measuring Cable burial insulates DAS from temperature transients, and temperature effects can generally be ignored at typical seismic and acoustic frequencies even on unburied cables.williams2023fiber
@@
In geophysics, DAS has been used for distributed dynamic temperature sensing to monitor fluid flow in boreholes.sidenko2022experimental ; titovspe ; karrenbach2019fiber ; bradley2024 Offshore DAS recordings in Japanide2021very , the Canary Islandswilliams2023fiber , and the Mediterranean Seapelaez2023high have revealed low-frequency transients over timescales from minutes to weeks, attributed to temperature oscillations from internal waves and turbulence. These are qualitatively consistent with observations from Raman-based DTS deployed on seafloor cablessinnett2020distributed ; davis2020fate and quantitatively consistent with collocated DTSgraeff2024shaken and DSTS measurementspelaez2023high in two cases.
@@ VII Conclusions
This review provides an encompassing description of DAS technology and demonstrates its potential applications to ocean acoustics. We detail how DAS converts a fiber-optic cable into a distributed sensor of vibrational fields, such as propagating sound, substantiating that active optic sensing can be used as a proxy for passive acoustic monitoring of the environment. Based on the mathematical derivation of the underlying physics from electromagnetic to mechanical and acoustic quantities in DAS measurements for coherent systems, we discuss the effect of typical DAS acquisition parameters in signal processing. The potential of DAS technology for underwater acoustic applications, such as sound source detection, is demonstrated with the Ocean Observatories Initiative data.
Deriving the directional sensitivity of the distributed sensing modality in terms of conventional array signal processing, we explain that a limitation of current DAS systems in observing high-frequency signals stems from competing data acquisition requirements. Specifically, the maximum detectable frequency is determined by the repetition frequency of the pulse that interrogates the optical fiber, which, in turn, is inversely proportional to the two-way travel time of the pulse along the fiber segment that defines the measurement range. Additionally, the longer the transmitted pulse length, the higher the signal-to-noise ratio of the measurand but the less sensitive the system becomes to high-frequency vibrational fields.
@@ Acknowledgements.
This work was performed while Angeliki Xenaki was employed at the NATO-STO Centre for Maritime Research and Experimentation and Peter Gerstoft was a visiting scientist. The work was funded by the NATO Allied Command Transformation and by the Office of Naval Research under Grant N00014-21-1-2267. The data supporting this study were collected by the Ocean Observatories Initiative (OOI), a major facility fully funded by the National Science Foundation under Cooperative Agreement No. 1743430, and the Woods Hole Oceanographic Institution OOI Program Office.
@@ Appendix A Coherent optical reflectometry
Optical time-domain reflectrometry detects the backscattered light from a pulse transmitted in an optical fiber with a photodetector in the interrogator [Hartog2017, , Sec. 3.1]. The photodetector’s output current $I_{pc}$ is proportional to the power of the received electric field, $I_{pc}\propto\lvert E_{b}\rvert^{2}$, hence the phase information of the backscattered signal is discarded. To retain the phase information of the backscttered light, coherent optical reflectrometry methods are used.
In principle, coherent optical reflectometry mixes the backscattered signal returning from the interrogated fiber $E_{b}$ with a reference signal $E_{\text{ref}}$. The resulting signal is fed to the photodetector such that the output current $I_{pc}$ is proportional to the power of the resulting interferometric field,He2021
@@
$=\lvert E_{\text{ref}}\rvert^{2}+\lvert E_{b}\rvert^{2}+2\Re\left\{E_{\text{ref}}E^{*}_{b}\right\}.$
@@
The reference signal $E_{\text{ref}}$ is, usually, a replica of the transmitted pulse $E_{p}$, potentially amplified and frequency shifted by an optical local oscillator,
$\widetilde{E}_{\text{ref}}(\tau)=G_{\text{LO}}e^{j\Delta\omega\tau}\widetilde{E}_{p}(\tau),$ (42)
@@
where $G_{\text{LO}}$ is the gain and $\Delta\omega$ the frequency shift introduced by the local oscillator. Deriving from Eqs. (42), (15), and (17), the interference product becomes,
$\widetilde{E}_{\text{ref}}(\tau)\widetilde{E}^{*}_{b}(\tau)$ (43)
@@
$=\quad G_{\text{LO}}e^{j\Delta\omega\tau}\widetilde{E}_{p}(\tau)\Big{(}\int\limits_{z-L_{p}/2}^{z+L_{p}/2}\!\!\widetilde{E}_{p}\left(\tau\!-\!2\frac{z_{s}}{c_{n}}\right)s(z_{s})\mathrm{d}z_{s}\Big{)}^{*}$
$=G_{\text{LO}}e^{j\Delta\omega\tau}p(\frac{\tau}{T_{p}})e^{j\omega_{p}\tau}$
$\quad\cdot\Big{(}\int\limits_{z-L_{p}/2}^{z+L_{p}/2}\!\!w(z-z_{s})\!e^{j\omega_{p}(\tau-2\frac{z_{s}}{c_{n}})}s(z_{s})\mathrm{d}z_{s}\Big{)}^{*}$
$=G_{\text{LO}}e^{j\left(\Delta\omega\tau+\phi(\tau)\right)}\int\limits_{z-L_{p}/2}^{z+L_{p}/2}\!\!w(z)w^{*}(z-z_{s})\!e^{j2k_{p}z_{s}}s^{*}(z_{s})\mathrm{d}z_{s}$
$=G_{\text{LO}}e^{j\left(\Delta\omega\tau+\phi(\tau)\right)}\widetilde{q}(z),$
where $\widetilde{q}(z)=\int_{z-L_{p}/2}^{z+L_{p}/2}w(z)w^{*}(z-z_{s})e^{j2k_{p}z_{s}}s(z_{s}^{*})\mathrm{d}z_{s}=q(z)e^{j\phi(z)}$ is introduced for notational brevity and $\phi(z)$ accounts for the phase introduced by changes in the optical path between the reference and the backscattered signal from location $z$ along the fiber received at time $\tau$.
@@
Hence, the photodetector output in Eq. (41) comprises the contributions of the signals $\widetilde{E}_{\text{ref}}$ at frequency $\omega_{p}+\Delta\omega$ [Eq. (42)], $\widetilde{E}_{b}$ at frequency $\omega_{p}$ [Eq. (17)], and the interference signal $\widetilde{E}_{\text{ref}}\widetilde{E}^{*}_{b}$ at the difference frequency $\Delta\omega$ [Eq. (43)]. Lowpass filtering the photodetector output $I_{pc}$ in Eq. (41) excludes the high-frequency contribution of the signals $|\widetilde{E}_{\text{ref}}|^{2}$ and $|\widetilde{E}_{b}|^{2}$ and retains the interference signal at the difference frequency $\Delta\omega$
$I_{\text{lp}}(\tau)$ $\propto\Re\left\{\widetilde{E}_{\text{ref}}(\tau)\widetilde{E}^{*}_{b}(\tau)\right\}$ (44)
$=\quad G_{\text{LO}}q(z)\cos\left(\Delta\omega\tau\!+\!\phi(z)\right)~.$
@@
Couplers can extract the phase of the detected signal [Hartog2017, , Sec. 3.3.1.1]. For example, an in-phase/quadrature (I/Q) demodulation systemwang2016coherent uses a $2\times 2$ channel coupler to output both the in-phase signal, Eq. (44), and the corresponding quadrature signal by phase shifting the in-phase signal by $\pi/2$,
$Q_{\text{lp}}(\tau)\!=\!I_{\text{lp}}(\tau,\pi/2)\!\propto\!G_{\text{LO}}q(z)\sin\left(\Delta\omega\tau+\phi(z)\right).$ (45)
Then, the phase of the detected signal is calculated as
$\phi(z)=\arctan\left(\frac{Q_{\text{lp}}(\tau)}{I_{\text{lp}}(\tau)}\right)-\Delta\omega\tau.$ (46)
By measuring the differential phase between the mixed signals, coherent optical reflectometry provides a measurand that is linearly related to the distance traveled by the backscattered wave.